# Any function I can actually write down is measurable, right?

### v2

@@ 3. Explicitly choosing things
Assuming $V=L$, for every basic Polish space $X$, there is a $\Delta^{1}_{2}$ well-ordering $<_{L}^{X}$ of $X$.
@@
If $X=\mathbb{R}$, Facts 2.10 and 3.1 and 2.11 establish that $<_{L}$ restricted to $\mathbb{R}^{2}$ is $\Sigma^{1}_{2}$. To see that it is also $\Pi^{1}_{2}$, note that $x\not<_{L}y$ if and only if $x=y$ or $y<_{L}x$. $x=y$ is a $\Pi^{0}_{1}$ (and therefore $\Sigma^{1}_{2}$) condition, so $x\not<_{L}y$ is $\Sigma^{1}_{2}$ by 2.7.
For other uncountable $X$, the result follows from Facts 2.9 and 3.1. ∎
Note that for powers of $\mathbb{R}$ in particular, we could also use the lexicographic order induced by $<_{L}$ restricted to $\mathbb{R}$. For example, we could take $\langle x,y\rangle<_{L}^{\mathbb{R}^{2}}\langle a,b\rangle$ to mean $x<_{L}a\vee(x=a\wedge y<_{L}b)$. The specifics do not matter too much.
@@
[…]
@@
[…]
@@
The fact that $[0,1]$ is $\Pi^{0}_{1}$ is obvious. It is also straightforward to show that the set $\{\langle x,x+q_{n},n\rangle:x\in\mathbb{R},~{}n\in\mathbb{N}\}\subseteq\mathbb{R}^{2}\times\mathbb{N}$ (for any given reasonable enumeration $(q_{n})_{n\in\mathbb{N}}$ of $\mathbb{Q}$) is $\Pi^{0}_{1}$, which implies that the set $\{\langle x,y\rangle:x-y\in\mathbb{Q}\}$ is $\Sigma^{0}_{2}$. By 2.7 and the fact that $[0,1]^{2}$ is also $\Pi^{0}_{1}$, this implies that $E$ is $\Sigma^{0}_{2}$. ∎
@@
By Lemmas 3.3 and 3.4, we can find a $\Pi^{1}_{2}$ Vitali set $A\subseteq[0,1]\subseteq\mathbb{R}$. By 2.11, there is a $\Pi^{1}_{1}$ $B\subseteq\mathbb{R}^{2}$ such that $A$ is the complement of $\pi_{2}(B)$. By 2.10, we can find a $\Pi^{0}_{2}$ set $C\subseteq\mathbb{R}^{2}\times\mathbb{N}$ such that $B$ is the complement of $\pi_{3}(C)$. By definition, there is a $\Pi^{0}_{2}$ set $D$ such that $C$ is the complement of $\pi_{4}(D)$. $U$ is then the complement of $D$. ∎
@@
$\mathsf{ZFC}$ is relatively consistent with the statement “every (lightface) $\Sigma^{1}_{n}$ subset of $\mathbb{R}$ is Lebesgue measurable for every $n$.”
@@
If all (boldface) $\undertilde{\mathbf{\Sigma}}^{1}_{3}$ sets of reals are Lebesgue measurable, then $\aleph_{1}$ is an inaccessible cardinal in $L$. In particular, the statement “all $\undertilde{\mathbf{\Sigma}}^{1}_{3}$ sets of reals are Lebesgue measurable” implies the consistency of $\mathsf{ZFC}$. (See Footnote 15.)
@@ 4.1. Hilbert’s tenth
Definition 4.2.
The original proof of the existence of universal polynomials, while constructive, didn’t do much to control the degree or number of unknowns of the resulting polynomial. There doesn’t even seem to be a published estimate of how big the original polynomial would be. Robinson and Matiyasevich, jointly and separately, produced a series of proofs reducing the needed number of unknowns, with Matiyasevich eventually reducing it to $9$ in an unpublished proof. Jones published this proof along with an analysis (some of which was performed by Wada) of various universal degree-unknown-number pairs, reporting the existence, for example, of a degree $4$ positively universal polynomial in $58$ unknowns and a $9$-unknown positively universal polynomial of degree $47216\cdot 5^{58}+9728\approx 1.638\times 10^{45}$ [11]. Unfortunately Jones omits a fair amount of details of these constructions and only explicitly gives the polynomial $p(x_{1},\dots,x_{28};n,m_{1},m_{2},m_{3})$ of degree $2\cdot 5^{60}$ shown in Figure 17 along with the general technique for constructing universal polynomials with $9$ variables. Focusing on this polynomial is reasonable, as it has the virtue of being short enough to actually write on one page, but it’s unclear how difficult it would be to fully replicated the other results reported by Jones. Using a standard technique of encoding intermediate calculations in extra variables [21, Thm. 6.2], it is possible to convert the polynomial in Figure 17 to an equivalent one of degree $4$ but with roughly $\log_{2}5^{60}\approx 140$ variables.
@@ 4.2. Polynomial engineering
For any $n$, any sequence of rational pairs $(a_{i},b_{i})_{1\leq i\leq n}$ with $a_{i}<b_{i}$ for each $i$, any real vector $\langle x_{1},\dots,x_{n}\rangle\in\prod_{i=1}^{n}(a_{i},b_{i})$, and any $s\in\mathbb{R}$, there are natural numbers $k_{1},\dots,k_{3+a}$ such that
@@
Let $r(\bar{x},\bar{k})=k_{1}^{2}k_{2}-k_{3}\sum_{i=1}^{n}(k_{1}x_{i}-k_{3+i}+k_{3})$. We clearly have that if $k_{1}$ and $k_{3}$ are positive, then the set of $\bar{x}$ for which $r(\bar{x},\bar{k})\geq 0$ is the closed Euclidean ball with center $\left\langle\frac{k_{4}-k_{3}}{k_{1}},\dots,\frac{k_{3+n}-k_{3}}{k_{1}}\right\rangle$ and radius $\sqrt{\frac{k_{2}}{k_{3}}}$. For any $\bar{x}\in\prod_{i=1}^{n}(a_{i},b_{i})$, we can find such a ball $B$ such that $B$ is a subset of $\prod_{i=1}^{n}(a_{i},b_{i})$ and $\bar{x}$ is in the interior of $B$. Moreover, we have that $r(\bar{x},t\bar{k})=t^{3}r(\bar{x},\bar{k})$. Therefore we can make the particular value of $r(\bar{x},\bar{k})$ arbitrarily large without changing the set $\{\bar{z}\in\mathbb{R}^{n}:r(\bar{z},\bar{k})\geq 0\}$. ∎
@@
$\sup_{\bar{y}\in\mathbb{N}^{\nu},\ell\in\mathbb{N},\bar{k}\in\mathbb{N}^{3+a}}p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=\begin{cases}1&n=0\\ +\infty&n>0,~{}\langle\bar{x},n-1\rangle\in U\\ 0&n>0,~{}\langle\bar{x},n-1\rangle\notin U\end{cases}$
@@
First note that clearly if $\ell_{1}\neq J(n,k_{1})$ or $\ell_{i+1}\neq J(\ell_{i},k_{i+1})$ for some $i\leq 2+a$, then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})\leq 0$. This is equivalent to saying that if $\ell_{3+a}\neq J_{4+a}(n,k_{1},k_{2},\dots,k_{3+a})$, then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})\leq 0$. Moreover if $q(\bar{y};\ell_{3+a},\bar{m})\neq 0$, then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})\leq 0$. Let $r(\bar{x},\bar{k})$ be the polynomial from the proof of 4.4. Note that if $\ell_{3+a}=J_{4+a}(n,\bar{k})$ and $q(\bar{y};\ell_{3+a},\bar{m})=0$, then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=r(\bar{x},\bar{k})$.
@@
  • $n>0$ and $k_{0}=k_{3}=0$, or
  • $n,k_{0},k_{3}>0$ and there is a $t\in W_{0}$ such that
@@
Note that this subset inclusion is decidable as a function of $\langle n,\bar{k}\rangle$ and $t$, so we have that $W_{1}$ is a computably enumerable set. Let $\bar{m}$ be a tuple of natural numbers chosen so that for any $\ell\in\mathbb{N}$, $q(\bar{y};\ell,\bar{m})$ has a solution in the natural numbers if and only if $\ell\in W_{1}$. Let $g(\bar{x},n)=\sup_{\bar{y}\in\mathbb{N}^{\nu},\bar{k}\in\mathbb{N}^{3+a}}p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})$.
Fix $\langle\bar{x},n\rangle\in\mathbb{R}^{a}\times\mathbb{N}$. If $n=1$, then we have that if $\ell_{3+a}=J_{4+a}(n,\bar{k})$, $k_{1}=2$, and $k_{2}=k_{3}=1$, then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=1$ and otherwise $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})\leq 0$. Therefore $g(\bar{x},n)=1$ whenever $n=1$.
If $n>1$ and $\ell_{3+a}=P_{4+a}(n,\bar{k})$ such that $k_{0}=k_{3}=1$, then we have that $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=0$. Therefore $g(\bar{x},n)\geq 0$. If $\langle\bar{x},n-1\rangle\in U$, we have by 4.4 that for any $s\in\mathbb{R}$, there are $\bar{y}$, $\bar{\ell}$, and $\bar{k}$ such that $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})>s$. Therefore $g(\bar{x},n)=+\infty$ for any $\langle\bar{x},n-1\rangle\in U$. If $\langle\bar{x},n-1\rangle\notin U$, then we have by construction that $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})\leq 0$ for any natural numbers $\bar{y}$, $\bar{\ell}$, and $\bar{k}$, so $g(\bar{x},n)\leq 0$ and therefore $g(\bar{x},n)=0$. ∎
@@
Fix a $\Pi^{0}_{2}$ set $X\subseteq\mathbb{R}^{a}$. By definition, there is a $\Sigma^{0}_{1}$ set $U\subseteq\mathbb{R}^{a}\times\mathbb{N}$ such that $X=\{\bar{x}\in\mathbb{R}^{a}:(\forall n\in\mathbb{N})\langle\bar{x},n\rangle\in U\}$. By 4.5, we can find a polynomial $p(x_{1},\dots,x_{a},n,k_{1},\dots,k_{6+\nu+2a})$ of the required degree such that $\sup_{\bar{k}\in\mathbb{N}^{6+\nu+2a}}p(\bar{x},n,\bar{k})$ is $1$ if $n=1$, +$\infty$+ ∞ if $n>1$ and $\langle\bar{x},n-1\rangle\in U$, and 00 if $n>1$ and $\langle\bar{x},n-1\rangle\notin U$. This implies that $\inf_{n\in\mathbb{N}}\sup_{\bar{k}\in\mathbb{N}^{6+\nu+2a}}p(\bar{x},n,\bar{k})$ is the indicator function of $X$. ∎
@@
By 3.5 there is a $\Pi^{1}_{2}$ set $A\subseteq\mathbb{R}$ such that $V=L$ implies $A$ is a Vitali subset of $[0,1]$. Applying 4.7 to the complement of $A$ (using the degree $4$ positively universal polynomial with $58$ unknowns in [11]) to get a polynomial $q$ and taking $1-q$ gives the required polynomial $p(x,y,z,n,k_{1},\dots,k_{70})$ of degree $7$. ∎
Of course, we can just as easily get that the indicator function of the well-ordering $<_{L}$ on $\mathbb{R}$ is similarly definable using a polynomial $p(x,y,z,w,n,k_{1},\dots,k_{72})$ of degree $7$ with integer coefficients using 3.1 and 4.7. Another thing we might want to try when contemplating variants of 4.8 is minimizing the number of variables required rather than the degree. We can attain the same result with a polynomial $p(x,y,z,n,k_{1},\dots,k_{14})$ of degree $973147216\cdot 5^{58}+$ 9731 using the suggested modification after 4.5 and the $9$-variable universal polynomial in [11].
It should be noted that $V=L$ is consistent with the presence of many (although not all) large cardinals, including the existence of a proper class of inaccessible cardinals (and more, such as Mahlo and some Erdős cardinals). This means that, on the one hand, the measurability of the function in 4.8 is independent of not just $\mathsf{ZFC}$, but of Tarski–Grothendieck set theory, [^19] Tarski–Grothendieck set theory is the theory $\mathsf{ZFC}+\text{``}$ “there is a proper class of inaccessible cardinals”, although the theory is not typically described this way specifically. which is strong enough to formalize the vast majority of proofs that occur outside of the context of set theory itself. On the other hand, the potential non-measurability of $f(x)$ occurs in extremely weak theories too, as low as even fragments of second-order arithmetic. This is because the construction of $L$ is robust enough to work even in such weak contexts [28, Sec. VII.4].
@@
has the same consistency strength (over $\mathsf{ZFC}$) as the existence of an inaccessible cardinal and in particular implies the consistency of $\mathsf{ZFC}$ and is therefore not provable in $\mathsf{ZFC}$.
@@
One direction is immediate from 4.10. The other direction (i.e., showing that all such function can consistently be measurable assuming the existence of an inaccessible cardinal) follows from [29, Thm. 2] and the fact that function $f(x)$ of the form in the given statement is $\undertilde{\mathbf{\Sigma}}^{1}_{n}$ for some $n$ (with $n$ at most one more than the number of infima and suprema in the definition of $f(x)$). ∎
@@
There is a polynomial $p(i,x,y,z,w,t,n,k_{1},\dots,k_{74})$ of degree $7$ with integer coefficients such that if $V=L$, then the functions $h(1,x,y,z),\dots,h(16,x,y,z)$ are the indicator functions of a paradoxical decomposition of $S^{2}$, where

### v3

@@ 3. Explicitly choosing things
Assuming $V=L$, for every basic Polish space $X$, there is a $\Delta^{1}_{2}$ well-ordering $_{L}^{X}$ of $X$.
@@
If $X=\mathbb{R}$, Facts 2.10 and 3.1 and 2.11 establish that $_{L}$ restricted to $\mathbb{R}^{2}$ is $\Sigma^{1}_{2}$. To see that it is also $\Pi^{1}_{2}$, note that $x\not<_{L}y$ if and only if $x=y$ or $y<_{L}x$. $x=y$ is a $\Pi^{0}_{1}$ (and therefore $\Sigma^{1}_{2}$) condition, so $x\not<_{L}y$ is $\Sigma^{1}_{2}$ by 2.7.
For other uncountable $X$, the result follows from 2.9. ∎
Note that for powers of $\mathbb{R}$ in particular, we could also use the lexicographic order induced by $_{L}$ restricted to $\mathbb{R}$. For example, we could take $\langle x,y\rangle<_{L}^{\mathbb{R}^{2}}\langle a,b\rangle$ to mean $x<_{L}a\vee(x=a\wedge y<_{L}b)$. The specifics do not matter too much.
@@
The set $\{\bar{x}\in\mathbb{R}^{n}:(\forall\bar{y}\in\mathbb{R}^{n})[\bar{x}\in B\wedge(\bar{y}\notin B\vee\bar{x}<_{L}^{\mathbb{R}^{n}}\bar{y}\vee\bar{x}=\bar{y}\vee\neg(\bar{x}\mathrel{E}\bar{y}))]\}$ defines a transversal of $E$ and is the complement of the projection of a $\Delta^{1}_{2}$ subset of $\mathbb{R}^{n}$ by 2.7. By 2.9, this implies that it is a $\Pi^{1}_{2}$ set. ∎
@@
$0,1]\subseteq\mathbb{R}$ is $\Pi^{0}_{1}$ and the equivalence relation $x\mathrel{E}y$ on $[0,1]$ defined by $x-y\in\mathbb{Q}$ is $\Sigma^{0}_{2}$. In particular, both are $\Delta^{1}_{2}$.
@@
The fact that $[0,1]$ is $\Pi^{0}_{1}$ is obvious. It is also straightforward to show that the set $\{\langle x,x+q_{n},n\rangle:x\in\mathbb{R},~n\in\mathbb{N}\}\subseteq\mathbb{R}^{2}\times\mathbb{N}$ (for any given reasonable enumeration $(q_{n})_{n\in\mathbb{N}}$ of $\mathbb{Q}$) is $\Pi^{0}_{1}$, which implies that the set $\{\langle x,y\rangle:x-y\in\mathbb{Q}\}$ is $\Sigma^{0}_{2}$. By 2.7 and the fact that $0,1]^{2}$ is also $\Pi^{0}_{1}$, this implies that $E$ is $\Sigma^{0}_{2}$. ∎
@@
By Lemmas 3.3 and 3.4, we can find a $\Pi^{1}_{2}$ Vitali set $A\subseteq[0,1]\subseteq\mathbb{R}$. By 2.11, there is a $\Pi^{1}_{1}$ $B\subseteq\mathbb{R}^{2}$ such that $A$ is the complement of $\pi_{2}(B)$. By 2.10, we can find a $\Pi^{0}_{2}$ set $C\subseteq\mathbb{R}^{3}$ such that $B$ is the complement of $\pi_{3}(C)$. By definition, there is a $\Sigma^{0}_{1}$ (i.e., semicomputable) set $U\subseteq\mathbb{R}^{3}\times\mathbb{N}$ such that $C$ is the complement of $\pi_{4}(\mathbb{R}^{3}\times\mathbb{N}\setminus U)$. So putting it all together we have that $A=\mathbb{R}\setminus\pi_{2}(B)=\mathbb{R}\setminus\pi_{2}(\mathbb{R}^{2}\setminus\pi_{3}(D))=\mathbb{R}\setminus\pi_{2}(\mathbb{R}^{2}\setminus\pi_{3}(\mathbb{R}^{3}\setminus\pi_{4}(\mathbb{R}^{3}\times\mathbb{N}\setminus U)))$. ∎
@@
$\mathsf{ZFC}$ is relatively consistent with the statement “every (lightface) $\Sigma^{1}_{n}$ subset of $\mathbb{R}$ is Lebesgue measurable for every nn.”
@@
If all (boldface) $\undertilde{\mathbf{\Sigma}}^{1}_{3}$ sets of reals are Lebesgue measurable, then $\aleph_{1}$ is an inaccessible cardinal in $L$. In particular, the statement “all $\undertilde{\mathbf{\Sigma}}^{1}_{3}$ sets of reals are Lebesgue measurable” implies the consistency of $\mathsf{ZFC}$. (See Figure 15.)
@@ 4.1. Hilbert’s tenth
Definition 4.2.
The original proof of the existence of universal polynomials, while constructive, didn’t do much to control the degree or number of unknowns of the resulting polynomial. There doesn’t even seem to be a published estimate of how big the original polynomial would be. Robinson and Matiyasevich, jointly and separately, produced a series of proofs reducing the needed number of unknowns, with Matiyasevich eventually reducing it to $9$ in an unpublished proof. Jones published this proof along with an analysis (some of which was performed by Wada) of various universal degree-unknown-number pairs, reporting the existence, for example, of a degree $4$ positively universal polynomial in $58$ unknowns and a $9$-unknown positively universal polynomial of degree $47216\cdot 5^{58}+9728\approx 1.638\times 10^{45}$ [14]. Unfortunately Jones omits a fair amount of details of these constructions and only explicitly gives the polynomial $p(x_{1},\dots,x_{28};n,m_{1},m_{2},m_{3})$ of degree $2\cdot 5^{60}$ shown in Figure 17 along with the general technique for constructing universal polynomials with $9$ variables. Focusing on this polynomial is reasonable, as it has the virtue of being short enough to actually write on one page, but it’s unclear how difficult it would be to fully replicate the other results reported by Jones. Using a standard technique of encoding intermediate calculations in extra variables [25, Thm. 6.2], it is possible to convert the polynomial in Figure 17 to an equivalent one of degree $4$ but with roughly $\log_{2}5^{60}\approx 140$ variables.
@@ 4.2. Polynomial engineering
For any $n$, any sequence of rational pairs $(a_{i},b_{i})_{1\leq i\leq n}$ with $a_{i}<b_{i}$ for each $i$, any real vector $\langle x_{1},\dots,x_{n}\rangle\in\prod_{i=1}^{n}(a_{i},b_{i})$, and any $s\in\mathbb{R}$, there are natural numbers $k_{1},\dots,k_{3+n}$ such that
@@
Let $r(\bar{x},\bar{k})=k_{1}^{2}k_{2}-k_{3}\sum_{i=1}^{n}(k_{1}x_{i}-k_{3+i}+k_{3})^{2}$. We clearly have that if $k_{1}$ and $k_{3}$ are positive, then the set of $\bar{x}$ for which $r(\bar{x},\bar{k})\geq 0$ is the closed Euclidean ball with center $\left\langle\frac{k_{4}-k_{3}}{k_{1}},\dots,\frac{k_{3+n}-k_{3}}{k_{1}}\right\rangle$ and radius $\sqrt{\frac{k_{2}}{k_{3}}}$. For any $\bar{x}\in\prod_{i=1}^{n}(a_{i},b_{i})$, we can find such a ball $B$ such that $B$ is a subset of $\prod_{i=1}^{n}(a_{i},b_{i})$ and $\bar{x}$ is in the interior of $B$. Moreover, we have that $r(\bar{x},t\bar{k})=t^{3}r(\bar{x},\bar{k})$. Therefore we can make the particular value of $r(\bar{x},\bar{k})$ arbitrarily large without changing the set $\{\bar{z}\in\mathbb{R}^{n}:r(\bar{z},\bar{k})\geq 0\}$. ∎
@@
$\sup_{\bar{y}\in\mathbb{N}^{\nu},\bar{\ell}\in\mathbb{N}^{3+a},\bar{k}\in\mathbb{N}^{3+a}}p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=\begin{cases}1&n=0\\ +\infty&n>0,~\langle\bar{x},n-1\rangle\in U\\ 0&n>0,~\langle\bar{x},n-1\rangle\notin U\end{cases}$
@@
  2. (ii)
    $n>0$ and $k_{1}=k_{3}=0$, or
  3. (iii)
    $n,k_{1},k_{3}>0$ and there is a $t\in W_{0}$ such that
@@
Note that the subset inclusion in (iii) is decidable as a function of $\langle n,\bar{k}\rangle$ and $t$, so we have that $W_{1}$ is a computably enumerable set. Let $\bar{m}$ be a tuple of natural numbers chosen so that for any $\ell\in\mathbb{N}$, $q(\bar{y};\ell,\bar{m})$ has a solution in the natural numbers if and only if $\ell\in W_{1}$.
Throughout the rest of this proof, we’ll say that the $\ell_{i}$’s are correct to mean that $\ell_{1}=J_{2}(n,k_{1})$ and $\ell_{i+1}=J_{2}(\ell_{i},k_{i+1})$ for all $i\leq 2+a$ (i.e., that the $\ell_{i}$’s have correctly coded the computation of $\ell_{3+a}=J_{4+a}(n,\bar{k})$). We’ll say that the $\ell_{i}$’s are not correct to mean that $\ell_{1}\neq J_{2}(n,k_{1})$ or $\ell_{i+1}=J_{2}(\ell_{i},k_{i+1})$ for some $i\leq 2+a$. Note that if the $\ell_{i}$’s are correct, then $(\ell_{1}-J_{2}(n,k_{1}))^{2}+\sum_{i=1}^{2+a}(\ell_{i+1}-J_{2}(\ell_{i},k_{i+1}))^{2}=0$ and if the $\ell_{i}$’s are not correct, then $(\ell_{1}-J_{2}(n,k_{1}))^{2}+\sum_{i=1}^{2+a}(\ell_{i+1}-J_{2}(\ell_{i},k_{i+1}))^{2}>0$.
Let $r(\bar{x},\bar{k})=k_{1}^{2}k_{2}-k_{3}\sum_{i=1}^{n}(k_{1}x_{i}-k_{3+i}+k_{3})^{2}$ be the polynomial from 4.4 again. Note that
$p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=r(\bar{x},\bar{k})-k_{1}^{2}k_{2}\left(q(\bar{y};\ell_{3+a},\bar{m})+(\ell_{1}-J_{2}(n,k_{1}))^{2}+\sum_{i=1}^{2+a}(\ell_{i+1}-J_{2}(\ell_{i},k_{i+1}))^{2}\right).$
We immediately have the following:
  • If $q(\bar{y};\ell_{3+a},\bar{m})=0$ and the $\ell_{i}$’s are correct, then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=r(\bar{x},\bar{k})$.
  • If $q(\bar{y};\ell_{3+a},\bar{m})>0$ or the $\ell_{i}$’s are not correct, then the first term of $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})$ is non-positive and so $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})\leq 0$.
Note that these two cases are exhaustive by the positive universality of $q(\bar{y};\ell_{3+a},\bar{m})$ and the choice of $\bar{m}$.
Let $g(\bar{x},n)=\sup_{\bar{y}\in\mathbb{N}^{\nu},\bar{\ell}\in\mathbb{N}^{3+a},\bar{k}\in\mathbb{N}^{3+a}}p(\bar{x},n,\bar{y},\bar{\ell},\bar{k}).$ We need to unpack the behavior of $g(\bar{x},n)$ to verify that it satisfies the conclusion of the lemma.
For $n=0$, we have the following:
  • If $k_{1}=1$, $k_{2}=k_{3}=0$ and the $\ell_{i}$’s are correct, then $\ell_{3+a}\in W_{1}$ by (i) and so $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=1$ for any choice of the other variables.
  • If $k_{1}\neq 1$, $k_{2}\neq 0$, $k_{3}\neq 0$, or the $\ell_{i}$’s are not correct, then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})\leq 0$ for any choice of the other variables.
Since these two cases are exhaustive, we have that $g(\bar{x},1)=1$ for any $\bar{x}$.
For $n>0$, we have the following:
  • If $k_{1}=k_{3}=0$ and the $\ell_{i}$’s are correct, then $\ell_{3+a}\in W_{1}$ by (ii) and so $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=0$ for any choice of the other variables.
  • If the conditions in (iii) are satisfied by $n$ and $\bar{k}$ and the $\ell_{i}$’s are correct, then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=r(\bar{x},\bar{k})$.
  • Otherwise if the conditions of the previous two bullet points fail, then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})\leq 0$.
The first bullet point implies that $g(\bar{x},n)\geq 0$ for any $\bar{x}$ and $n>0$. The definition of $W_{1}$ implies that for any $\bar{x}\notin U$, if $n>0$ and $\bar{k}$ satisfy (iii), then $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=r(\bar{x},\bar{k})\leq 0$. So for any $\bar{x}$ and $n>0$ with $\langle\bar{x},n-1\rangle\notin U$, we have that $g(\bar{x},n)=0$. Finally, for any $\bar{x}$ and $n>0$ with $\langle\bar{x},n-1\rangle\in U$, we have by 4.4 that for any $s\in\mathbb{R}$, there are $\bar{y}$, $\bar{\ell}$, and $\bar{k}$ such that $p(\bar{x},n,\bar{y},\bar{\ell},\bar{k})=r(\bar{x},\bar{k})>s$. Therefore for any $\langle\bar{x},n-1\rangle\in U$, $g(\bar{x},n)=+\infty$. ∎
@@
Fix a $\Pi^{0}_{2}$ set $X\subseteq\mathbb{R}^{a}$. By definition, there is a $\Sigma^{0}_{1}$ set $U\subseteq\mathbb{R}^{a}\times\mathbb{N}$ such that $X=\{\bar{x}\in\mathbb{R}^{a}:(\forall n\in\mathbb{N})\langle\bar{x},n\rangle\in U\}$. By 4.5, we can find a polynomial $p(x_{1},\dots,x_{a},n,k_{1},\dots,k_{6+\nu+2a})$ of the required degree such that $\sup_{\bar{k}\in\mathbb{N}^{6+\nu+2a}}p(\bar{x},n,\bar{k})$ is $1$ if $n=1$, $\infty$ if $n>1$ and $\langle\bar{x},n-1\rangle\in U$, and 0 if $n>1$ and $\langle\bar{x},n-1\rangle\notin U$. This implies that $\inf_{n\in\mathbb{N}}\sup_{\bar{k}\in\mathbb{N}^{6+\nu+2a}}p(\bar{x},n,\bar{k})$ is the indicator function of $X$. ∎
@@
By 3.5 there is a $\Pi^{1}_{2}$ set $A\subseteq\mathbb{R}$ such that $V=L$ implies AA is a Vitali subset of $[0,1]$. Applying 4.7 to the complement of AA (using the degree $4$ positively universal polynomial with $58$ unknowns in [14]) to get a polynomial $q$ and taking 1q1-q gives the required polynomial $p(x,y,z,n,k_{1},\dots,k_{70})$ of degree $7$. ∎
Of course, we can just as easily get that the indicator function of the well-ordering $_{L}$ on $\mathbb{R}$ is similarly definable using a polynomial $p(x,y,z,w,n,k_{1},\dots,k_{72})$ of degree $7$ with integer coefficients using 3.1 and 4.7. Another thing we might want to try when contemplating variants of 4.8 is minimizing the number of variables required rather than the degree. We can attain the same result with a polynomial $p(x,y,z,n,k_{1},\dots,k_{14})$ of degree $47216\cdot 5^{58}+9731$ using the suggested modification after 4.5 and the $9$-variable universal polynomial in [14].
It should be noted that $V=L$ is consistent with the presence of many (although not all) large cardinals, including the existence of a proper class of inaccessible cardinals (and more, such as Mahlo and some Erdős cardinals). This means that, on the one hand, the measurability of the function in 4.8 is independent of not just $\mathsf{ZFC}$, but of Tarski–Grothendieck set theory, [^19] Tarski–Grothendieck set theory is the theory $\mathsf{ZFC}+\text{``}$there is a proper class of inaccessible cardinals”, although the theory is not typically described this way specifically. which is strong enough to formalize the vast majority of proofs that occur outside of the context of set theory itself. On the other hand, the potential non-measurability of $f(x)$ occurs in extremely weak theories too, as low as even fragments of second-order arithmetic. This is because the construction of $L$ is robust enough to work even in such weak contexts [32, Sec. VII.4].
@@
has the same consistency strength (over $\mathsf{ZFC}$) as the existence of an inaccessible cardinal. In particular, it implies the consistency of $\mathsf{ZFC}$ and is therefore not provable in $\mathsf{ZFC}$.
@@
One direction is immediate from 4.10. The other direction (i.e., showing that all such function can consistently be measurable assuming the existence of an inaccessible cardinal) follows from [33, Thm. 2] and the fact that function $f(x)$ of the form in the given statement is $\undertilde{\mathbf{\Sigma}}^{1}_{n}$ for some $n$ (with $n$ at most one more than the number of infima and suprema in the definition of $f(x)$). ∎
@@
There is a polynomial $p(m,x,y,z,w,t,n,k_{1},\dots,k_{76})$ of degree $7$ with integer coefficients such that if $V=L$, then the functions $h(1,x,y,z),\dots,h(16,x,y,z)$ are the indicator functions of a paradoxical decomposition of $S^{2}$, where